Find the velocity of a particle if it describes 10 feet in 2 second This question is of calculus chapter

Grade:10

2 Answers

Arun
4 years ago
Dear student
 
As velocity is a vector quantity, hence curve is required.
For speed = 10/2 = 5 feet per second
Vikas TU
4 years ago
Dear stuent 
 
Let us assume the directon of feet is upward direction , 
So , j will be the direction of velocity . 
10 feet in 2 second 
V = 10/2 j 
V = 5j 
Hope this helps 
Good Luck